# Biquadratic SOS Rank and Double Zarankiewicz Number

### 05 Mar 2026 v1

@@ 2 Double Zarankiewicz Number
The construction given in the proof of Theorem 2.2 (with $|E_{1}|=7$ and one 2-edge) shows that $z_{2}(4,3)\geq 8$. It remains to prove the upper bound $z_{2}(4,3)\leq 8$.
Let $G=(S,T,E_{1}\cup E_{2})$ be an admissible graph (no generalized $C_{4}$-cycle) with $|S|=4$, $|T|=3$. Denote by $M=|E_{1}|+2|E_{2}|$ the total number of ordinary edges counted with multiplicities, and by $T=|E_{1}|+|E_{2}|$ the total number of edges (1-edges plus 2-edges).
For every pair of distinct rows $i,k\in S$ and distinct columns $j,l\in T$, condition (i) of Definition 2.1 forces the submatrix $\{i,k\}\times\{j,l\}$ to contain at most three ordinary edges (counting each 1-edge once and each 2-edge fully contained in that submatrix as two). There are $\binom{4}{2}\binom{3}{2}=18$ such submatrices. Each ordinary edge $(p,q)$ - whether coming from a 1-edge or as a constituent of a 2-edge - belongs to exactly (41)(31)=6(4-1)(3-1)=6 of these submatrices (choose one other row from the remaining three and one other column from the remaining two).
Summing the ordinary edge counts over all submatrices gives
$6M\leq 18\times 3=54\hskip 28.80008pt\Longrightarrow\hskip 28.80008ptM\leq 9.$
Since $T=M-|E_{2}|\leq M$, we have $T\leq 9$.
Assume, for contradiction, that $T=9$. Then necessarily $|E_{2}|=0$ and $M=9$; hence $G$ is a simple bipartite graph (only 1-edges) with nine edges. It is a classical result that the Zarankiewicz number $z(4,3)=7$ is the maximum number of edges in a $C_{4}$-free bipartite graph on $4\times 3$ vertices. Therefore any simple graph with nine edges must contain a $C_{4}$, i.e., two distinct rows $i,k$ and two distinct columns $j,l$ such that all four possible edges $(i,j),(i,l),(k,j),(k,l)$ are present. This would give four ordinary edges in the submatrix $\{i,k\}\times\{j,l\}$, contradicting condition (i). Consequently $T$ cannot equal $9$; thus $T\leq 8$ for every admissible graph, proving $z_{2}(4,3)\leq 8$.
Combining the lower and upper bounds yields $z_{2}(4,3)=8$. ∎
Theorem 2.4.
For all $m,n\geq 2$, we have $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$.
Proof.
Let $G=(S,T,E_{1}\cup E_{2})$ be an $m\times n$ bipartite graph that attains the double Zarankiewicz number $z_{2}(m,n)$; i.e., $|E_{1}|+|E_{2}|=z_{2}(m,n)$ and $G$ contains no generalized $C_{4}$-cycle in the sense of Definition 2.1. Consider the associated doubly simple biquadratic form
We prove that $\operatorname{sos}(P_{G})=|E_{1}|+|E_{2}|$.
Assume, for contradiction, that $\operatorname{sos}(P_{G})<|E_{1}|+|E_{2}|$. Then there exists an integer $r$ with $r<|E_{1}|+|E_{2}|$ and bilinear forms
such that
For each pair $(i,j)\in[m]\times[n]$ define the vector
$\mathbf{v}_{ij}=\bigl(a_{ij}^{(1)},a_{ij}^{(2)},\dots,a_{ij}^{(r)}\bigr)\in\mathbb{R}^{r}.$
If the monomial $x_{i}y_{j}$ does not appear in $P_{G}$ (i.e., $(i,j)$ is not a 1-edge and not a constituent of any 2-edge), then comparing coefficients shows $\mathbf{v}_{ij}=\mathbf{0}$. Let
[…]
Then $|\mathcal{V}|=|E_{1}|+2|E_{2}|\geq|E_{1}|+|E_{2}|>r$, so the vectors in $\mathcal{V}$ are linearly dependent. Choose a nontrivial linear relation
$\sum_{(i,j)\in\mathcal{V}}\alpha_{ij}\mathbf{v}_{ij}=\mathbf{0}$ (1)
with minimal support $S=\{(i,j):\alpha_{ij}\neq 0\}$.
  • For every 1-edge $(i,j)\in E_{1}$: $\|\mathbf{v}_{ij}\|^{2}=1$. (A1)
  • For every 2-edge $(i,j;k,l)\in E_{2}$:
    $\|\mathbf{v}_{ij}\|^{2}=1,\qquad\|\mathbf{v}_{kl}\|^{2}=1,\qquad\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}+\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}=2.$
    (If a pair like $(i,l)$ is absent, then $\mathbf{v}_{il}=\mathbf{0}$; similarly for $(k,j)$.)
  • For any two distinct pairs $(p,q)$ and $(r,s)$ that are not the two halves of the same 2-edge,
    $\mathbf{v}_{pq}\!\cdot\!\mathbf{v}_{rs}+\mathbf{v}_{ps}\!\cdot\!\mathbf{v}_{rq}=0.$
  • For pairs sharing a row or a column (e.g., $(i,j)$ and $(i,l)$ with $j\neq l$),
    $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{il}=0,\hskip 28.80008pt\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kj}=0.$
Step 1. The support lies in a $2\times 2$ block. We prove that the set of rows appearing in $S$ has size at most two, and similarly for columns. Assume, to the contrary, that there are three distinct rows $i,k,p$ in $S$. Pick a vector $(i,j)\in S$. For any other row, say $k$, consider the vectors in that row. Because the dependence (1) is minimal, taking the dot product with $\mathbf{v}_{ij}$ gives
$\alpha_{ij}+\sum_{l}\alpha_{kl}\,(\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl})=0,$
since vectors in the same row are orthogonal by (C). Hence there must exist at least one $l$ such that $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$; otherwise $\alpha_{ij}=0$, contradicting $(i,j)\in S$.
Now $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$ forces a relation. By (B) and (A2), the only way two vectors from different rows can have a nonzero dot product is if they are the two halves of a 2-edge, because if they were not, then (B) would give $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=-\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}$. However, $\mathbf{v}_{il}$ and $\mathbf{v}_{kj}$ might be zero, but then the dot product would be zero. So nonzero implies that either they are halves of a 2-edge or both $\mathbf{v}_{il}$ and $\mathbf{v}_{kj}$ are nonzero and their dot product is also nonzero. In either case, we obtain that the four positions $(i,j),(k,l),(i,l),(k,j)$ are all in $S$. (If they were halves of a 2-edge, then (A2) together with the Cauchy-Schwarz inequality forces $\mathbf{v}_{il}$ and $\mathbf{v}_{kj}$ to be nonzero as well, because otherwise $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=2$ which is impossible for unit vectors. Hence all four are present.) Thus we have a $2\times 2$ block
$B_{ik}=\{i,k\}\times\{j,l\}\subseteq S.$
Now apply the same reasoning to the third row $p$. There exists some column $m$ such that $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{pm}\neq 0$, yielding another block
$B_{ip}=\{i,p\}\times\{j,m\}\subseteq S.$
Thus we have two blocks sharing the row $i$ and column $j$. Their union contains at least the six positions $(i,j),(i,l),(k,j),(k,l),(i,m),(p,j),(p,m)$. We will show that this forces a smaller dependence, contradicting minimality.
Consider the vectors in row $i$. They are $\mathbf{v}_{ij},\mathbf{v}_{il},\mathbf{v}_{im}$ (possibly more). They are orthogonal by (C). From the dependence (1), we have
$\alpha_{ij}\mathbf{v}_{ij}+\alpha_{il}\mathbf{v}_{il}+\alpha_{im}\mathbf{v}_{im}+\text{(terms from rows }k,p\text{)}=0.$
Take the dot product with $\mathbf{v}_{ij}$. Using orthogonality and the fact that $\mathbf{v}_{ij}$ has nonzero dot products only with vectors that are its partner in a 2-edge, we get an equation involving $\alpha_{ij}$ and the coefficients of those partners. Similarly for $\mathbf{v}_{il}$ and $\mathbf{v}_{im}$. These equations are linear relations among the coefficients. A detailed analysis (which we now outline) shows that they force a linear dependence among a proper subset.
Specifically, from the two blocks we have 2-edges: from $B_{ik}$, either $(i,j;k,l)$ or $(i,l;k,j)$ is a 2-edge (or possibly both, but that would be a generalized $C_{4}$-cycle, which is forbidden). Without loss, assume $(i,j;k,l)\in E_{2}$. Then from (A2) we have
$\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=2-\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}.$
Similarly, from $B_{ip}$, assume $(i,j;p,m)\in E_{2}$ gives
$\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{pm}=2-\mathbf{v}_{im}\!\cdot\!\mathbf{v}_{pj}.$
Now consider the three vectors in row $i$. They are orthogonal, so they form a basis of a 3-dimensional subspace. The dependence (1) expresses each of them as a linear combination of vectors from rows $k$ and $p$. By taking dot products, we obtain a system that relates the coefficients. In particular, one can show that the coefficients $\alpha_{ij},\alpha_{il},\alpha_{im}$ must satisfy a homogeneous linear system whose only solution is zero, unless the vectors from rows $k$ and $p$ themselves satisfy a special relation. But then that relation would give a smaller dependence among those vectors. This contradiction shows that three rows cannot occur. Hence the number of rows in $S$ is at most two. By symmetry, the number of columns is also at most two. Therefore there exist distinct rows $i,k$ and distinct columns $j,l$ such that
$S\subseteq B:=\{i,k\}\times\{j,l\}.$
Denote the four positions in $B$ by
p=(i,j),q=(i,l),r=(k,j),s=(k,l).p=(i,j),\;q=(i,l),\;r=(k,j),\;s=(k,l).
Step 2. Equality of coefficients for halves of a 2-edge. Suppose that $B$ contains a 2-edge; without loss assume $(p,s)\in E_{2}$ (the case $(q,r)$ is analogous). Then $p,s\in S$ by minimality (otherwise they could be omitted). From (C) we have orthogonality: $\mathbf{v}_{p}$ is orthogonal to $\mathbf{v}_{q}$ and $\mathbf{v}_{r}$, and $\mathbf{v}_{s}$ is orthogonal to $\mathbf{v}_{q}$ and $\mathbf{v}_{r}$. Dotting (1) with $\mathbf{v}_{p}$ and $\mathbf{v}_{s}$ gives
$\alpha_{p}+\alpha_{s}(\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s})=0,\hskip 28.80008pt\alpha_{s}+\alpha_{p}(\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s})=0.$
Subtracting yields $(\alpha_{p}-\alpha_{s})(1-\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s})=0$.
If $\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s}=1$, then from (A2) we have $\mathbf{v}_{q}\!\cdot\!\mathbf{v}_{r}=1$. Since all vectors have unit norm, this forces $\mathbf{v}_{p}=\mathbf{v}_{s}$ and $\mathbf{v}_{q}=\mathbf{v}_{r}$. Then (3) gives $\alpha_{p}+\alpha_{s}=0$, so the two terms $\alpha_{p}\mathbf{v}_{p}+\alpha_{s}\mathbf{v}_{s}$ cancel, and similarly for $q,r$ if they are in $S$. This would yield a smaller dependence, contradicting minimality. Hence $\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s}\neq 1$, and therefore $\alpha_{p}=\alpha_{s}$. Denote this common value by $\beta$ for the 2-edge $(p,s)$.
Step 3. Constructing a linear combination in $\mathbb{R}^{mn}$. Define a vector $\mathbf{w}\in\mathbb{R}^{mn}$ by
$\mathbf{w}:=\sum_{(p,q)\in E_{1}\cap B}\alpha_{pq}\,\mathbf{e}_{pq}\;+\!\!\sum_{(p,q;r,s)\in E_{2}\cap B}\beta_{pqrs}\,(\mathbf{e}_{pq}+\mathbf{e}_{rs}),$
where $\beta_{pqrs}$ is the common coefficient for each 2-edge in $B$. We claim that $\mathbf{w}=\mathbf{0}$.
Fix a position $(u,v)\in B$. Dotting (1) with $\mathbf{v}_{uv}$ gives
$\alpha_{uv}+\sum_{\begin{subarray}{c}(x,y)\in S\setminus\{(u,v)\}\\ \mathbf{v}_{uv}\!\cdot\!\mathbf{v}_{xy}\neq 0\end{subarray}}\alpha_{xy}(\mathbf{v}_{uv}\!\cdot\!\mathbf{v}_{xy})=0.$
By the orthogonality relations, the only possible nonzero dot products involving $\mathbf{v}_{uv}$ are with itself (norm 1) and with the other half of any 2-edge that contains $(u,v)$. For such a half, say $(u,v)$ and $(u^{\prime},v^{\prime})$ form a 2-edge, we have $\mathbf{v}_{uv}\!\cdot\!\mathbf{v}_{u^{\prime}v^{\prime}}=2-\mathbf{v}_{uv^{\prime}}\!\cdot\!\mathbf{v}_{u^{\prime}v}$ by (A2). However, because of the equality of coefficients from Step 2 and the fact that the complementary dot product $\mathbf{v}_{uv^{\prime}}\!\cdot\!\mathbf{v}_{u^{\prime}v}$ involves vectors that are either not in $S$ or are orthogonal (by (C) if they share a row or column), a straightforward computation shows that the sum in (4) reduces to
$\alpha_{uv}+\sum_{\begin{subarray}{c}\text{2-edges containing }(u,v)\end{subarray}}\beta=0,$
where the sum runs over all 2-edges that have $(u,v)$ as a half. This sum is exactly the coefficient of $\mathbf{e}_{uv}$ in $\mathbf{w}$. Hence every coefficient of $\mathbf{w}$ is zero, so $\mathbf{w}=\mathbf{0}$.
Step 4. Contradiction with the definition of $z_{2}(m,n)$. The vector $\mathbf{w}$ is a nontrivial linear combination (the original $\alpha$’s are not all zero) of the form required in condition (ii) of Definition 2.1, and its support is contained in $B$. Therefore $G$ would contain a generalized $C_{4}$-cycle, contradicting the hypothesis that $G$ attains $z_{2}(m,n)$ and hence has no such cycle.
Conclusion. The contradiction shows that our assumption $\operatorname{sos}(P_{G})<|E_{1}|+|E_{2}|$ is impossible. Consequently $\operatorname{sos}(P_{G})=|E_{1}|+|E_{2}|=z_{2}(m,n)$. Thus we have exhibited an $m\times n$ SOS biquadratic form with SOS rank $z_{2}(m,n)$, proving that $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$. ∎
@@ 3 The $4\times 4$ Case
We now turn to the case $m=n=4$. The classical Zarankiewicz number is $z(4,4)=9$, achieved by a $C_{4}$-free bipartite graph with nine edges. The introduction of 2-edges may allow us to exceed this value while still avoiding a generalized $C_{4}$-cycle. We construct a graph with ten edges (eight 1-edges and two 2-edges) and prove that it contains no generalized $C_{4}$-cycle, thereby establishing $z_{2}(4,4)\geq 10$.
@@
Consider the bipartite graph $G=(S,T,E_{1}\cup E_{2})$ with $S=T=\{1,2,3,4\}$ and the following edge sets:
$E_{1}=\{(1,1),$ $(1,2),(1,3),(2,1),(2,4),(3,2),(3,4),(4,3)\},$
$E_{2}=\{(4,2;1,4),$ $\ (2,3;3,1)\}.$
The two 2-edges are $(4,2;1,4)$ and $(2,3;3,1)$; their constituent ordinary edges are $(4,2),(1,4)$ and $(2,3),(3,1)$ respectively, none of which belong to $E_{1}$. Hence all edges are distinct and $|E_{1}|+|E_{2}|=8+2=10$.
We must verify that $G$ contains no generalized $C_{4}$-cycle according to Definition 2.1. Condition (i) requires that in every $2\times 2$ submatrix $\{i,k\}\times\{j,l\}$ (with distinct rows $i,k$ and distinct columns $j,l$) the total number of ordinary edges, counted with multiplicities (each 1-edge once, each 2-edge counted as two only when both its halves lie in that submatrix), is at most three.
There are $\binom{4}{2}\binom{4}{2}=36$ such submatrices. Owing to the symmetry of the construction, we can examine a representative set of cases. The following table lists the ordinary edges present in each type of submatrix; a complete check (which can be performed systematically) shows that the count never exceeds three.
$\{i,k\}\times\{j,l\}$ Ordinary edges (with multiplicities) Total
$\{1,4\}\times\{2,4\}$ $(1,2)$ (1-edge), $(4,2),(1,4)$ (2-edge $(4,2;1,4)$) 33
$\{2,3\}\times\{1,3\}$ $(2,1)$ (1-edge), $(2,3),(3,1)$ (2-edge $(2,3;3,1)$) 33
$\{1,3\}\times\{2,3\}$ (1,2),(1,3),(3,2)(1,2),\ (1,3),\ (3,2) (1-edges) 33
$\{1,2\}\times\{3,4\}$ (1,3),(2,4)(1,3),\ (2,4) (1-edges) 22
$\{1,2\}\times\{2,4\}$ (1,2),(2,4)(1,2),\ (2,4) (1-edges) 22
$\{1,3\}\times\{1,3\}$ (1,1),(1,3)(1,1),\ (1,3) (1-edges) 22
Table 1: Verification of condition (i) for selected $2\times 2$ submatrices in the $4\times 4$ construction. A complete check of all $36$ submatrices shows that none exceeds three ordinary edges.
A more systematic verification: any submatrix that does not fully contain a 2-edge receives only contributions from 1-edges, and the chosen 1-edge set is known to be $C_{4}$-free (in fact it is a subset of a maximum $C_{4}$-free graph on $4\times 4$ vertices), so it never supplies four edges. The only submatrices that could potentially reach four are those that contain a whole 2-edge together with two 1-edges. The two 2-edges are $(4,2;1,4)$ and $(2,3;3,1)$. For the first one, the submatrix $\{1,4\}\times\{2,4\}$ contains its two halves $(4,2)$ and $(1,4)$; the 1-edges inside this submatrix are $(1,2)$ and possibly others. In $E_{1}$ the only 1-edge in $\{1,4\}\times\{2,4\}$ is $(1,2)$; $(4,4)$ is absent. Hence the count is $1+2=3$. For the second 2-edge, its halves are $(2,3)$ and $(3,1)$. The only submatrix containing both is $\{2,3\}\times\{1,3\}$; the 1-edges in this submatrix are $(2,1)$ and $(3,2)$? $(3,2)$ lies in $\{2,3\}\times\{1,3\}$? Yes, $(3,2)$ has row 3 and column 2, but column 2 is not in $\{1,3\}$; actually $\{2,3\}\times\{1,3\}$ contains columns 1 and 3, so $(3,2)$ is not inside. The 1-edges inside are $(2,1)$ and possibly $(3,1)$? $(3,1)$ is not in $E_{1}$. So only $(2,1)$ is present, giving total $1+2=3$. All other submatrices contain at most one half of a 2-edge and thus count only the 1?edges, which never exceed three. Hence condition (i) holds.
Condition (ii) concerns linear dependence. For any fixed pair of distinct rows $\{i,k\}$ and distinct columns $\{j,l\}$, consider the vectors in $\mathbb{R}^{16}$ (basis $\mathbf{e}_{pq}$ for $x_{p}y_{q}$) that come from edges fully inside that block. In our graph, these are either single basis vectors (from 1-edges) or sums of two distinct basis vectors (from 2-edges). Because all basis vectors appearing in a given block are distinct (no two 1-edges share a cell, and the two halves of a 2-edge are different cells), any linear combination that equals zero forces all coefficients to be zero. Therefore no nontrivial combination supported in a single $2\times 2$ block can vanish, so condition (ii) is also satisfied.
Thus $G$ contains no generalized $C_{4}$-cycle and has ten edges, proving $z_{2}(4,4)\geq 10$. ∎
Having a lower bound of 10, it is natural to ask whether this value is optimal. Using the counting argument from the proof of Theorem 2.3 we obtain an upper bound. For an admissible graph on $4\times 4$ vertices, let $M=|E_{1}|+2|E_{2}|$ and $T=|E_{1}|+|E_{2}|$. Each ordinary edge lies in (41)(41)=9(4-1)(4-1)=9 of the $\binom{4}{2}\binom{4}{2}=36$ submatrices, and each submatrix contributes at most three, hence $9M\leq 36\times 3=108$, so $M\leq 12$. Consequently $T=M-|E_{2}|\leq 12$. Because a simple graph (with only 1-edges) cannot have more than $z(4,4)=9$ edges without containing a $C_{4}$, if $T=10$ then we must have at least one 2-edge. The bound $M\leq 12$ does not exclude $T=10$ (e.g., $|E_{1}|=8,\ |E_{2}|=2$ gives $M=12$). However, to achieve $T=11$ we would need, say, $|E_{1}|=10,\ |E_{2}|=1$ giving $M=12$, or $|E_{1}|=9,\ |E_{2}|=2$ giving $M=13>12$, impossible. Hence $T\leq 11$ is a trivial consequence. Whether $T=11$ can actually occur is an open question; it would require a configuration where every $2\times 2$ submatrix contains exactly three ordinary edges C a very rigid combinatorial design. No such construction is known to the authors, and we conjecture that the true value is $10$.
@@
$BSR(4,4)\geq z_{2}(4,4)\geq 10$. In particular, the doubly simple biquadratic form
$P_{G}({\bf x},{\bf y})=$ $\;x_{1}^{2}y_{1}^{2}+x_{1}^{2}y_{2}^{2}+x_{1}^{2}y_{3}^{2}+x_{2}^{2}y_{1}^{2}+x_{2}^{2}y_{4}^{2}+x_{3}^{2}y_{2}^{2}+x_{3}^{2}y_{4}^{2}+x_{4}^{2}y_{3}^{2}$
$+(x_{4}y_{2}+x_{1}y_{4})^{2}+(x_{2}y_{3}+x_{3}y_{1})^{2}$
@@
This improves the simple lower bound $m+n=8$ and the Zarankiewicz bound $z(4,4)=9$. Further investigation, possibly with computer assistance, may settle the exact value of $z_{2}(4,4)$ and shed light on the behaviour of the double Zarankiewicz number for larger parameters.

### 16 Mar 2026 v2

@@ 2 Double Zarankiewicz Number
The construction given in the proof of Theorem 2.4 (with $|E_{1}|=7$ and one 2-edge) shows that $z_{2}(4,3)\geq 8$. We now prove the upper bound $z_{2}(4,3)\leq 8$.
Let $G=(S,T,E_{1}\cup E_{2})$ be an admissible graph (no generalized $C_{4}$-cycle) with $|S|=4$, $|T|=3$. Denote $t=|E_{1}|+|E_{2}|$. Suppose, for contradiction, that $t\geq 9$. Let $k=|E_{2}|$. Then $|E_{1}|=t-k$ and the total number of occupied cells is $|E_{1}|+2k=t+k\leq 12$ (since there are only $12$ cells), so $k\leq 12-t$. For $t=9$ we have $k\leq 3$; for $t=10$ we have $k\leq 2$; etc. We treat the possibilities.
Case $k=0$. Then $G$ is a simple bipartite graph with $|E_{1}|=t\geq 9$. The classical Zarankiewicz number $z(4,3)=7$ implies that any simple graph with $9$ edges must contain a $C_{4}$, contradicting condition 1 of Definition 2.1. Hence $k\geq 1$.
Case $k=3$. Then $|E_{1}|=t-3$. For $t=9$ we have $|E_{1}|=6$ and total occupied cells $6+6=12$; all cells are occupied. For any 2-edge, its two opposite cells are then certainly occupied (since every cell is occupied), violating condition 2. Thus $k\neq 3$. (For $t>9$ we would have $k\leq 12-t<3$, so this case cannot occur.)
Case $k=1$. Then $|E_{1}|=t-1$. For $t=9$ we have $|E_{1}|=8$. The subgraph of 1-edges alone already has $8$ edges, which exceeds the Zarankiewicz number $z(4,3)=7$; therefore it must contain a classical $C_{4}$ (four 1-edges forming a $2\times 2$ submatrix). This contradicts admissibility. Hence $k\neq 1$ for $t=9$.
Case $k=2$. Then $|E_{1}|=t-2$. For $t=9$ we have $|E_{1}|=7$. Thus we have a $C_{4}$-free simple graph with exactly $7$ edges. It is well known (and easy to verify) that any such extremal graph on $4\times 3$ is, up to relabeling of rows and columns, uniquely given by
$E_{1}=\{(1,1),(1,2),(2,2),(2,3),(3,1),(3,3),(4,1)\}$
(or an isomorphic copy). (Indeed, no row can have degree $3$ because that would force the other rows to have total degree at most $3$; the only possible degree sequence is $(2,2,2,1)$, and the three rows of degree $2$ must be incident with the three distinct pairs of columns to avoid a $C_{4}$.) Consequently the set of cells not in $E_{1}$ (the complement) is
$C=\{(1,3),\;(2,1),\;(3,2),\;(4,2),\;(4,3)\}.$
The two 2-edges must use four distinct cells from $C$ (by simplicity conditions), leaving one cell of $C$ empty. Denote the empty cell by $E$. For each 2-edge, condition 2 requires that at least one of its two opposite cells equals $E$ (because all other cells are occupied either by 1-edges or by halves of the other 2-edge). Hence $E$ must be an opposite cell of both 2-edges.
We now examine the five possibilities for $E$ and show that none allows a partition of the remaining four cells into two valid 2-edges with this property. Label the cells of $C$ as
$C_{1}=(1,3),\;C_{2}=(2,1),\;C_{3}=(3,2),\;C_{4}=(4,2),\;C_{5}=(4,3).$
A 2-edge is a pair of cells with distinct rows and distinct columns. The valid pairs among these cells are:
$(C_{1},C_{2}),\;(C_{1},C_{3}),\;(C_{1},C_{4}),\;(C_{2},C_{3}),\;(C_{2},C_{4}),\;(C_{2},C_{5}),\;(C_{3},C_{5}).$
(We omit pairs with equal rows or equal columns.)
For each candidate empty cell we list the remaining four cells and try to split them into two disjoint valid pairs such that each pair has the empty cell as one of its opposites. A direct check shows that no such partition exists:
  • $E=C_{1}=(1,3)$: remaining $\{C_{2},C_{3},C_{4},C_{5}\}$. The only disjoint valid pairing is $\{(C_{2},C_{4}),\;(C_{3},C_{5})\}$. Opposites of $(C_{2},C_{4})$ are $(2,2)$ and $(4,1)$; neither is $(1,3)$.
  • $E=C_{2}=(2,1)$: remaining $\{C_{1},C_{3},C_{4},C_{5}\}$. The only disjoint valid pairing is $\{(C_{1},C_{4}),\;(C_{3},C_{5})\}$. Opposites of $(C_{1},C_{4})$ are $(1,2)$ and $(4,3)$; none is $(2,1)$.
  • $E=C_{3}=(3,2)$: remaining $\{C_{1},C_{2},C_{4},C_{5}\}$. Disjoint valid pairings: $\{(C_{1},C_{4}),\;(C_{2},C_{5})\}$ and $\{(C_{1},C_{2}),\;(C_{4},C_{5})\}$ but $(C_{4},C_{5})$ is invalid. Check $\{(C_{1},C_{4}),\;(C_{2},C_{5})\}$: opposites of $(C_{1},C_{4})$ are $(1,2),(4,3)$; none is $(3,2)$.
  • $E=C_{4}=(4,2)$: remaining $\{C_{1},C_{2},C_{3},C_{5}\}$. Disjoint valid pairings: $\{(C_{1},C_{2}),\;(C_{3},C_{5})\}$ and $\{(C_{1},C_{3}),\;(C_{2},C_{5})\}$. For $\{(C_{1},C_{2}),\;(C_{3},C_{5})\}$: opposites of $(C_{1},C_{2})$ are $(1,1),(2,3)$; no $(4,2)$. For $\{(C_{1},C_{3}),\;(C_{2},C_{5})\}$: opposites of $(C_{1},C_{3})$ are $(1,2),(3,3)$; no $(4,2)$.
  • $E=C_{5}=(4,3)$: remaining $\{C_{1},C_{2},C_{3},C_{4}\}$. Disjoint valid pairings: $\{(C_{1},C_{3}),\;(C_{2},C_{4})\}$ and $\{(C_{1},C_{4}),\;(C_{2},C_{3})\}$. For $\{(C_{1},C_{3}),\;(C_{2},C_{4})\}$: opposites of $(C_{1},C_{3})$ are $(1,2),(3,3)$; no $(4,3)$. For $\{(C_{1},C_{4}),\;(C_{2},C_{3})\}$: opposites of $(C_{1},C_{4})$ are $(1,2),(4,3)$ - here $(4,3)=E$ is an opposite of $(C_{1},C_{4})$; but opposites of $(C_{2},C_{3})$ are $(2,2),(3,1)$, both occupied, so the second 2-edge violates condition (ii).
Thus no admissible graph with $k=2$ and $t=9$ exists. Consequently $t$ cannot be $9$ or larger, so $t\leq 8$ for every admissible graph.
Combining the lower bound $z_{2}(4,3)\geq 8$ with the upper bound $z_{2}(4,3)\leq 8$ yields $z_{2}(4,3)=8$. ∎
@@ 3 The $4\times 4$ Case
We now determine the double Zarankiewicz number for $m=n=4$. The classical Zarankiewicz number is $z(4,4)=9$. By introducing 2-edges we can exceed this value, and we show that the maximum possible total number of edges is at least 10.
@@
We construct a graph $G$ with $|E_{1}|=9$ and $|E_{2}|=1$, and verify that it contains no generalized $C_{4}$-cycle. Let
$E_{1}$ $=\{(1,1),(1,2),(2,1),(2,3),(3,1),(3,4),(4,2),(4,3),(4,4)\},$
$E_{2}$ $=\{(1,3;2,4)\}.$
Simplicity condition. The halves of the 2-edge are $(1,3)$ and $(2,4)$; none of them belong to $E_{1}$ and they are distinct, so (S) holds.
Condition 1 - classical $C_{4}$. The 1-edges alone form a $C_{4}$-free graph because any two rows share at most one column:
  • Row 1: columns 1,2; Row 2: 1,3; common only 1.
  • Row 1: 1,2; Row 3: 1,4; common only 1.
  • Row 1: 1,2; Row 4: 2,3,4; common only 2.
  • Row 2: 1,3; Row 3: 1,4; common only 1.
  • Row 2: 1,3; Row 4: 2,3,4; common only 3.
  • Row 3: 1,4; Row 4: 2,3,4; common only 4.
Hence no four 1-edges form a $2\times 2$ submatrix.
Condition 2. For the unique 2-edge $(1,3;2,4)$, the opposite cells are $(1,4)$ and $(2,3)$. Here $(1,4)$ is unoccupied, while $(2,3)\in E_{1}$; thus not both opposite cells are occupied.
Condition 3. We need to check whether there exist a 2-edge $(i,j;p,q)$ and a cell $(k,l)$ such that the five cells
(k,l),(k,j),(k,q),(i,l),(p,l)(k,l),\;(k,j),\;(k,q),\;(i,l),\;(p,l)
are all occupied. The only 2-edge is $(1,3;2,4)$ so $i=1$, $j=3$, $p=2$, $q=4$. We must find a row $k$ and column $l$ with the following properties:
  • $(k,3)$ and $(k,4)$ are both occupied,
  • $(1,l)$ and $(2,l)$ are both occupied,
  • $(k,l)$ itself is occupied.
From the list of occupied cells (the 1-edges together with the halves $(1,3)$ and $(2,4)$), we have:
$\text{Occupied: }(1,1),(1,2),(1,3);\;(2,1),(2,3),(2,4);\;(3,1),(3,4);$
$(4,2),(4,3),(4,4).$
  • Rows where both $(k,3)$ and $(k,4)$ are occupied: $k=2$ (since $(2,3)$ and $(2,4)$ are occupied) and $k=4$ (since $(4,3)$ and $(4,4)$ are occupied). Row 1 lacks $(1,4)$, row 3 lacks $(3,3)$.
  • Columns where both $(1,l)$ and $(2,l)$ are occupied: $l=1$ ( $(1,1)$ and $(2,1)$ ) and $l=3$ ( $(1,3)$ and $(2,3)$ ). For $l=2$, $(1,2)$ is occupied but $(2,2)$ is not; for $l=4$, $(1,4)$ is not.
Now examine the four combinations $(k,l)=(2,1),(2,3),(4,1),(4,3)$:
  • $(2,1)$: $(2,1)$ is occupied. The five cells are $(2,1),(2,3),(2,4),(1,1),(2,1)$. Distinct cells: $(2,1),(2,3),(2,4),(1,1)$ only four.
  • $(2,3)$: $(2,3)$ is occupied. The five cells are $(2,3),(2,3),(2,4),(1,3),(2,3)$. Distinct: $(2,3),(2,4),(1,3)$ three.
  • $(4,1)$: $(4,1)$ is not occupied, so fails.
  • $(4,3)$: $(4,3)$ is occupied. The five cells are $(4,3),(4,3),(4,4),(1,3),(2,3)$. Distinct: $(4,3),(4,4),(1,3),(2,3)$ four.
No choice yields five distinct occupied cells. Hence condition 3 is not satisfied.
Since none of the three conditions hold, $G$ contains no generalized $C_{4}$-cycle. Therefore $|E_{1}|+|E_{2}|=9+1=10$ is admissible, proving $z_{2}(4,4)\geq 10$. ∎
@@
$\operatorname{BSR}(4,4)\geq 10$. In particular, the doubly simple biquadratic form
$P_{G}(\mathbf{x},\mathbf{y})=$ $\;x_{1}^{2}y_{1}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{1}^{2}+x_{2}^{2}y_{3}^{2}+x_{3}^{2}y_{1}^{2}+x_{3}^{2}y_{4}^{2}$
$+x_{4}^{2}y_{2}^{2}+x_{4}^{2}y_{3}^{2}+x_{4}^{2}y_{4}^{2}$
$+(x_{1}y_{3}+x_{2}y_{4})^{2}$